optimal choice

If P1=P2, then at the optimal choice

x1 must be equal to x2
X and Y must be normal goods
MU1 may equal MU2 but it is not necessarily so
MU1 must equal MU2

One way to look at the consumer's optimal choice is as one in which the marginal utility of the last dollar spent

is higher in the more preferred goods
is equal to -P1/P2
in all goods is the same
is equal to the marginal utility per dollar saved.

If MU1/P1 < MU2/P2 and the consumer is already spending all of his income, then which of the following is correct

the consumer gets a higher satisfaction per dollar spent on x1
the consumer get a higher satisfaction per dollar spent on x2
the consumer gets a higher satisfaction from the last dollar spent on x1
the consumer gets a higher satisfaction from the last dollar spent on x2
Answer #1

Answer to Q1:

MU1 must equal MU2

Optimal bundle is that bundle where MU1/MU2=P1/P2. So,if p1=p2,then MU1= MU2.

Answer to Q2:

in all goods is the same.

Ans.3- the consumer gets a higher satisfaction from the last dollar spent on x2 . This is because if MU1/P1<MU2/P2,then our marginal utility for the last dollar spent is higher on good 2,so consumer gets higher satisfaction from good 2.



answered by: Zahidul Hossain